# From “The Cliff” to “Virgil”: Mapping the Spectral Diversity of Little Red Dots with JWST/NIRSpec

### v1

@@ Abstract
One of JWST’s most unexpected discoveries is the emergence of “Little Red Dots” (LRDs): compact sources at with blue rest-frame UV continua, red optical slopes, and broad Balmer emission lines that challenge standard models and suggest a population of early, unusual active galactic nuclei (AGNs). Using a comprehensive photometric selection and public NIRSpec/PRISM spectroscopy across six JWST deep fields, we identify a large sample of 118 LRDs with high-S/N spectra, enabling a population-wide analysis of their UV–optical continuum and emission lines. We find clear correlations between rest-frame color ([0.30.9 m]) and slopes: bluer LRDs have blue UV slopes () and red optical slopes, while redder LRDs exhibit redder UV slopes (). The continuum shape shows a similar trend: redder LRDs display prominent Balmer breaks and curvature, while bluer LRDs follow power-law–like optical SEDs. From literature compilations, 60% of known broad-line AGNs satisfy our LRD criteria, and up to 90% of LRDs show broad Balmer lines. Emission-line diagnostics reveal a shift from high H /H and low [O iii]5007/H in redder LRDs to the opposite in bluer ones, along with stronger narrow-line equivalent widths, suggesting a transition from AGN- to host-dominated emission. We fit the spectra with a two-component model combining a gas-enshrouded black hole (BH) and a galaxy host. Redder LRDs require higher-luminosity, unreddened BHs and modestly reddened hosts; bluer LRDs require lower-luminosity, reddened BHs and dust-free galaxies. This framework reproduces the diversity in colors and spectral shape by varying BH luminosity, obscuration, and host–to-BH luminosity ratio.
1 Introduction
@@
On the spectroscopic front, JWST/NIRSpec has delivered over 200 spectra of LRDs, revealing emission lines in nearly all cases and firmly establishing the widespread presence of broad-line emission across the population (greene23; kocevski24; taylor24; juodzbalis25; hviding25). Just as critical, continuum detections have significantly shaped physical interpretations of LRDs. Initial reports of Balmer breaks in several sources pointed to stellar-dominated continua (wang24_lrd; wang24b; labbe24b). However, further observations revealed larger breaks increasingly inconsistent with stellar populations (labbe24b; ma25; ji25), culminating in extreme cases with break strengths above 5 magnitudes (degraaff25; naidu25; taylor25) that challenged the stellar origin.
@@
In this paper, we address these issues by constructing a large, uniformly selected sample of photometric LRDs across six JWST deep fields. Using an inclusive, color-based selection, we recover the full photometric diversity seen in earlier studies, including sources missed by more restrictive criteria. We then cross-match this sample with all available NIRSpec spectroscopy compiled in the Dawn JWST Archive (DJA)111https://dawn-cph.github.io/dja , enabling a systematic comparison of photometric and spectroscopic properties. We use continuum spectroscopy to measure UV and optical slopes free from emission line contamination, allowing us to characterize color gradients, broad-line incidence, and continuum shapes. This unified approach provides a consistent baseline for evaluating which spectroscopic features are representative of the broader LRD population.
@@
2 Data
2.1 JWST/NIRCam Photometry
For source identification and photometry in the CEERS-EGS, PRIMER-UDS, PRIMER-COSMOS, and UNCOVER fields, we use version v7.2 of the publicly available photometric catalogs from the DJA. These catalogs are derived from JWST/NIRCam Level-2 imaging processed with the grizli pipeline (brammer_grizli), including PSF-matching across filters. Source detection is performed using SEP (SEP), a Python implementation of SExtractor (sex), applied to a detection image typically based on F444W. Photometry is measured in 05 diameter circular apertures and corrected to total flux using Kron scaling (kron80), with filter-dependent aperture corrections applied based on empirical PSFs (see valentino23 and Brammer et al., in prep., for details). For the JADES fields (GOODS-N and GOODS-S), we use the official Data Release 2 (DR2) multi-band photometric catalogs (eisenstein24), which include consistent photometry across NIRCam bands based on the JADES pipeline.
2.2 JWST/NIRSpec Spectroscopy
@@
3 Photometric selection of LRDs
3.1 Selection criteria and completeness
@@
Overall, we identify 1000 LRDs down to a limiting magnitude of F444W (28 in the JADES fields), including 157 LRDs with NIRSpec/PRISM spectra in DJA. Approximately 70% of the spectra have sufficient signal-to-noise in the rest-frame UV and optical continuum for SED fitting (see § 4 and 5). Sample sizes by field are summarized in Table 1. Figure 1 shows the LRD color–color selection diagram from barro25 for all six fields, with the photometric LRD population shown as an orange density map and spectroscopic LRDs as gray circles. Most spectroscopically identified LRDs from the literature fall within our photometric selection region, demonstrating excellent agreement between photometric and spectroscopic classifications. This includes all NIRCam/WFSS BLAGNs from matthee24 and 76% of the red, high–equivalent-width (EW,Å) sources from setton24. The remaining 20% are predominantly extended sources (setton24) with red rest-frame UV–optical SEDs characteristic of dusty, massive galaxies (e.g., 6585-COS-60504, RUBIES-UDS-41598, RUBIES-UDS-152282, RUBIES-UDS-19735, WIDE-GDN-6406). Excluding these extended sources increases the selection completeness to 87%.
@@
There are, however, a small number of compact, spectroscopic LRDs with broad emission lines that are missed by our photometric selection, as well as by most other criteria. These sources typically have relatively low redshifts () and/or unusually red rest-frame UV slopes, which shift their F115WF200W colors out of the LRD selection region and into the locus of massive dusty galaxies (see Figure 1 of barro24). We identify seven such objects in the literature and include them manually in the photometric sample for completeness; they are marked with yellow circles in Figure 1. Notably, RUBIES-BLAGN-1 (wang24b), JADES-1180-12402 (setton24; williams23), the “Rosetta” AGN (juodzbalis24a), and RUBIES-UDS-144195 (hviding25) are among the lowest-redshift LRDs, at , 3.19, 2.26, and 3.404, respectively. All but the “Rosetta” source show significantly redder UV slopes (see discussion in § 4). We additionally include the three reported extreme Balmer break sources (“black hole stars”, BH⋆) from degraaff25, naidu25, and taylor25, which exhibit weak UV emission and are identified solely by their extremely red optical colors.
@@
While the addition of these six spectroscopic LRDs contributes negligibly to the overall photometric sample, and less than 1% to the spectroscopic one, they highlight the need for a more complete diagnostic criterion based on rest-frame continuum properties. Classification using rest-frame UV and optical slopes (kocevski24; hainline24) or UV-to-optical colors (barro24) offers a more robust way to organize and distinguish LRD subtypes. As we show in § 4, these six sources fall at the red end of the rest-frame color diagram but are otherwise consistent with the broader LRD population.
@@
3.2 Other compact, non-LRD sources
Compact quiescent galaxies represent another population that can overlap with and contaminate LRD selections. These sources share key observable features, red colors, compact morphologies, and, in some cases, broad emission lines (Kocevski et al. 2024), but their intrinsic SEDs differ markedly from those of LRDs. Specifically, quiescent galaxies exhibit red UV slopes, and while their optical continua are steep near the Balmer break, they flatten rapidly at longer wavelengths, deviating from the approximate power-law behavior typical of LRDs.
@@ 3.2 Other compact, non-LRD sources
| Field | Phot. LRDs | NIRSpec LRDs | S/N | % BLAGNs LRDs | % LRDs BL |
| CEERS-EGS | 127 | 45 | 25 | 61% | 89% |
| PRIMER-UDS | 353 | 62 | 33 | 59% | 89% |
| JADES-GDS | 92 | 18 | 17 | 50% | 70% |
| JADES-GDN | 112 | 18 | 16 | 67% | 54% |
| UNCOVER | 45 | 15 | 12 | 92% | 65% |
| PRIMER-COSMOS | 207 | 23 | 15 | N/A | N/A |
| Total | 968 | 157 | 118 | - | - |
3.3 Spectroscopic samples of BLAGNs
@@
Overall, we find that the fraction of LRDs among samples of BLAGNs is relatively high, typically ranging from 60% up to 90% in UNCOVER. These values are higher than those reported in earlier studies using the same BLAGN samples, primarily due to (1) the adoption of a more flexible photometric selection that extends to bluer UV-to-optical colors (i.e., shallower optical slopes), and (2) the inclusion of a small number of spectroscopic LRDs at lower redshift that fall outside traditional color cuts. For example, taylor24 report an LRD fraction of 34% among BLAGN in CEERS-EGS and PRIMER-UDS, while hainline24 find a value of 30% in JADES after excluding three LRD-targeted sources. Indeed, the spectroscopic follow-up strategy plays a significant role in these differences. The 90% fraction in UNCOVER reflects a strong targeting bias, as many sources were selected specifically to follow up the LRDs identified in labbe23. Similarly, the RUBIES survey prioritized red NIRCam sources, which naturally favors LRD-like objects. As noted by greene23, both the equivalent width and FWHM of broad lines tend to be larger in the reddest sources (e.g., F277WF444W 1.6), making BLAGNs that are also LRDs easier to detect. hainline24 reach a similar conclusion, showing that LRDs dominate the high-luminosity end of the AGN population. Consequently, it is clear that larger, unbiased spectroscopic samples are needed to reliably assess the intrinsic fraction of LRDs among the full BLAGN population.
The fraction of LRDs with broad emission lines is also relatively high, typically around 50–90%. Importantly, these values represent lower limits, as nearly all LRDs show clear emission lines, and the identification of broad components is often limited by the depth and sensitivity of the observations. However, the presence of broad lines alone does not necessarily imply that the AGN dominates the continuum in any given spectral range.
4 Rest-frame UV slopes and optical colors
@@
4.1 Rest-frame color - slope diagram classification
Here, we analyze a subset of 118 LRDs (out of 157) with NIRSpec/PRISM spectroscopy and high S/N continuum in the UV. We require an average S/N per pixel in the range Å. Since LRDs are typically red, this threshold also implies higher S/N in the optical, ensuring sufficient signal for reliable SED modeling. We compute UV slopes by fitting a power law to the continuum between the Lyman and Balmer limits, using the convention . The slope is measured in rather than the traditional , where , to enable direct comparison with the rest-frame UV-to-optical color, [0.30.9 m] , defined using synthetic tophat filters of width 100 Å.
@@ 4.1 Rest-frame color - slope diagram classification
This trend poses a challenge for photometric selections, which are tuned to detect the combination of blue UV and red optical slopes. Instead, the reddest LRDs begin to resemble quiescent or dusty galaxies with uniformly red continua, making them harder to distinguish in standard color–color diagrams (as noted in Section 3.1). Several LRDs with red UV slopes were thus identified only a posteriori from spectroscopy (wang24_lrd; juodzbalis24a; williams24). Yet once placed on the color–slope diagram, they fall well within the main LRD distribution, supporting their classification as true LRDs.
@@
4.2 Curvature vs. power-law in the rest-optical SED
We use the detailed spectral coverage of the optical SED provided by NIRSpec/PRISM to examine the continuum shape of LRDs. Our goal is to quantify what fraction of the population shows significant curvature or breaks in the optical, which may point to stellar-dominated SEDs (e.g., williams24; pg24a), or more exotic BH⋆ or quasi-stellar components (e.g., inayoshi24). These cases differ from the smoother power-law continua often associated with canonical obscured AGNs in broadband photometry (e.g., polletta06).
@@
4.3 The variety of SED types in LRDs
@@
The reddest bin ([0.30.9 m] mag) contains LRDs with the strongest curvature, reddest UV slopes, and most pronounced continuum breaks, with an average Balmer break strength of 4.41, far beyond what standard stellar population models can reproduce. This bin includes the three extreme LRDs from degraaff25, naidu25, and taylor25 that motivated the black hole star (BH⋆) model, in which a dense gas cocoon surrounds an accreting black hole, causing the sharp decline. We further discuss this modeling in § 5. In addition to their breaks, these LRDs exhibit sharp flattening toward the NIR, consistent with the weak MIRI colors reported for the reddest LRDs (e.g., degraaff25; setton25; ronayne25). The adjacent bin ( [0.30.9 m] mag) shows similar overall shapes, with somewhat weaker curvature, red UV slopes, and a lower average Balmer break amplitude of 2.35. This group includes some of the first spectroscopically confirmed LRDs with strong but less extreme Balmer breaks (kocevski24; labbe24b; wang24b). In both of the two reddest bins, LRDs tend to exhibit prominent, broad Balmer emission lines, with weaker or undetected metal lines. We discuss line ratios and equivalent widths in more detail in § 4.7.
The three bluer bins correspond to LRDs with [0.30.9 m] , approaching the median color of the photometric sample ([0.30.9 m] mag). These sources display a remarkably similar flat UV continuum, with , much bluer than in the redder bins. As [0.30.9 m] decreases from 3 to 1, the average Balmer break amplitude declines steadily to 1.57, 1.13, and 1.07, values more typical of normal blue star-forming galaxies. Similarly, SED curvature diminishes toward shallower, power-law–like optical slopes. In these bins, the composite SED closely resembles the canonical combination of blue plus red power laws, with an inflection near the Balmer limit. In parallel, the median spectra show a growing prominence of narrow metal lines, especially the coronal lines around the Balmer break, such as [O ii] and [Ne iii], while the relative strength of H appears to decline compared to the blended H +[O iii]5007 complex.
Notably, the median SEDs in the two bluest bins suggest that the inflection may occur redward of H∞, near the [O iii]5007 emission line. This is consistent with “Virgil” (iani24; rinaldi25), a blue LRD with a seemingly flat optical continuum that rises sharply beyond H . A shift in the inflection is quite relevant. One of the main arguments against two-component models (e.g., galaxy+AGN) has been the consistency of the inflection at H∞, which would favor a single-component origin (setton24). If the inflection point varies with color, it would instead favor a multi-component scenario for the peculiar SED of LRDs. We revisit this scenario in detail in § 5.
4.4 Breakers vs. Power-Laws: Diversity and Degeneracy in LRD SEDs
@@
The left panel shows LRDs with pronounced curvature, following the trend between [0.30.9 m] color and curvature strength (; see color-coding in Figure 3). The reddest, highest-curvature sources exhibit the most extreme Balmer breaks: their SEDs rise steeply in the optical and then flatten sharply toward the NIR. Moving toward bluer LRDs, the break becomes less pronounced, the optical slope shallower, and the flattening shifts to longer wavelengths. A simple interpretation in the context of a two-component model, would be that the relative luminosity of the break-dominated component gradually decreases top-to-bottom, effectively hiding the steep slope and pushing the inflection point redward.
The right panel shows power-law LRDs. While most are blue, a small fraction (10%) exhibit very red colors (Figure 3). These red power-law sources show a range of optical slopes, suggesting variation in the physical conditions shaping the continuum, such as dust attenuation or gas density in BH⋆ models. Notably, the reddest of these LRDs, JADES-GDN-68797 and MACSJ0647.7+7015 (one of the first reported LRDs; killi23), exhibit remarkably similar optical slopes, despite the latter being about 1 mag bluer. This also agrees with the apparent shift in the inflection point from the flat UV to the rising optical continuum in this object, which occurs at longer wavelengths. A similar shift is seen in “Virgil” (yellow) and other bluer LRDs, as discussed in the previous section. We return to this point in § 5.
4.5 Color Dependence of the Emission Line FWHMs

### v2

@@ Abstract
One of JWST’s most unexpected discoveries is the emergence of “Little Red Dots” (LRDs): compact sources at with blue rest-frame UV continua, red optical slopes, and broad Balmer emission lines that challenge standard models and suggest a population of early, unusual active galactic nuclei (AGNs). Using a comprehensive photometric selection and public NIRSpec/PRISM spectroscopy across six JWST deep fields, we identify a large sample of 118 LRDs with high-S/N spectra, enabling a population-wide analysis of their UV–optical continuum and emission lines. We find clear correlations between rest-frame UV–optical color ([0.30.9 m]) and slopes: bluer LRDs have blue UV slopes () and red optical slopes, while redder LRDs exhibit redder UV slopes (). The continuum shape shows a similar trend: redder LRDs display prominent Balmer breaks and curvature, while bluer LRDs follow power-law–like optical SEDs. From literature compilations, 60% of known broad-line AGNs satisfy our LRD criteria, and up to 90% of LRDs show broad Balmer lines. Emission-line diagnostics reveal a shift from high H /H and low [O iii]5007/H in redder LRDs to the opposite in bluer ones, along with stronger narrow-line equivalent widths, suggesting a transition from AGN- to host-dominated emission. We fit the spectra with a two-component model combining a gas-enshrouded black hole (BH) and a galaxy host. Redder LRDs require higher-luminosity, unreddened BHs and modestly reddened hosts; bluer LRDs require lower-luminosity, reddened BHs and dust-free galaxies. This framework reproduces the diversity in colors and spectral shape by varying BH luminosity, obscuration, and BH-to-host luminosity ratio.
I Introduction
@@
On the spectroscopic front, JWST/NIRSpec MSA (ferruit22) has delivered over 200 spectra of LRDs, revealing emission lines in nearly all cases and firmly establishing the widespread presence of broad-line emission across the population (greene23; kocevski24; taylor24; juodzbalis25; hviding25). Just as critical, continuum detections have significantly shaped physical interpretations of LRDs. Initial reports of Balmer breaks in several sources pointed to stellar-dominated continua (wang24_lrd; wang24b; labbe24b). However, further observations revealed larger breaks increasingly inconsistent with stellar populations (labbe24b; ma25; ji25), culminating in extreme cases with break strengths above 5 magnitudes (degraaff25; naidu25; taylor25) that challenged the stellar origin.
@@
In this paper, we address these issues by constructing a large, uniformly selected sample of photometric LRDs across six JWST deep fields. In § III, we apply an inclusive, color-based selection that recovers the full photometric diversity reported in previous studies, including sources missed by more restrictive criteria. We then cross-match this sample with all available NIRSpec spectroscopy compiled in the Dawn JWST Archive (DJA)111https://dawn-cph.github.io/dja, enabling a systematic comparison between photometric and spectroscopic properties. In § IV, we use continuum spectroscopy to measure rest-frame UV and optical slopes and colors, free from emission-line contamination, and to characterize population-wide correlations among these quantities. Finally, in § V, we interpret these trends within a semi-empirical two-component model combining a galaxy and a BH⋆ component, demonstrating that it naturally reproduces the observed range of LRD continuum and emission-line properties.
@@
II Data
II.1 JWST/NIRCam Photometry
For source identification and photometry in the CEERS-EGS, PRIMER-UDS, PRIMER-COSMOS, and UNCOVER fields, we use version v7.2 of the publicly available photometric catalogs from the DJA. These catalogs are derived from JWST/NIRCam Level-2 imaging processed with the grizli pipeline (brammer_grizli), including PSF-matching across filters. Source detection is performed using SEP (SEP), a Python implementation of SExtractor (sex), applied to a detection image typically based on F444W. Photometry is measured in 05 diameter circular apertures and corrected to total flux using Kron scaling (kron80), with filter-dependent aperture corrections applied based on empirical PSFs (see valentino23 and Brammer et al., in prep., for details). For the JADES fields (GOODS-N and GOODS-S), we use the official Data Release 3 (DR3) multi-band photometric catalogs (eisenstein24; deugeniojades25), which include consistent photometry across NIRCam bands based on the JADES pipeline.
II.2 JWST/NIRSpec Spectroscopy
@@
III Photometric selection of LRDs
III.1 Selection criteria and completeness
@@
Overall, we identify 1000 LRDs down to a limiting magnitude of F444W (28 in the JADES fields), including 191 LRDs with NIRSpec/PRISM spectra in DJA over an area of 632.4 arcmin2. Approximately 70% of the spectra have sufficient signal-to-noise in the rest-frame UV and optical continuum for SED fitting (see § IV and V). Sample sizes by field are summarized in Table 1. Figure 1 shows the redshift distribution for the full photometric LRD sample (gray) and the NIRSpec/PRISM spectroscopic subset (black, scaled by a factor of 3). The photometric sample spans a wide range from to 9, with two prominent peaks at and , and a median redshift of . This is consistent with barro25, but our analysis expands the sample from 248 to nearly 1000 LRDs and extends coverage to all six JWST deep fields. As in barro25, the lower-redshift peak is partly driven by LRDs with bluer optical colors and strong [O iii]5007 emission that boosts F277W; these are recovered by our selection and by kokorev24, but are largely missed by redder cuts such as those in kocevski24, which yield a higher median redshift of . In contrast, our selection becomes increasingly incomplete at , where the F150W-F200W color begins tracing the Lyman break. The spectroscopic subset shows a slightly lower median redshift of , likely reflecting selection biases toward brighter targets more easily observed in follow-up. This is consistent with their brighter median F444W, compared to 26.2 for the full photometric sample.
@@
Figure 2 shows the LRD color–color selection diagram from barro25 for all six fields, with the photometric LRD population shown as an orange density map and spectroscopic LRDs as gray circles. Most spectroscopically identified LRDs from the literature fall within our photometric selection region, demonstrating excellent agreement between photometric and spectroscopic classifications. This includes all NIRCam/WFSS BLAGNs from matthee24 and 76% of the red, high–equivalent-width (EW,Å) sources from setton24. The remaining 20% are predominantly extended sources (setton24) with red rest-frame UV–optical SEDs characteristic of dusty, massive galaxies (e.g., 6585-COS-60504, RUBIES-UDS-41598, RUBIES-UDS-152282, RUBIES-UDS-19735, WIDE-GDN-6406). Excluding these extended sources increases the selection completeness to 87%.
@@
There are, however, a small number of compact, spectroscopic LRDs with broad emission lines that are missed by our photometric selection, as well as by most other criteria. These sources typically have relatively low redshifts () and/or unusually red rest-frame UV slopes, which shift their F115WF200W colors out of the LRD selection region and into the locus of massive dusty galaxies (see Figure 1 of barro25). We identify seven such objects in the literature and include them manually in the photometric sample for completeness; they are marked with yellow circles in Figure 2. Notably, RUBIES-BLAGN-1 (wang24b), JADES-1180-12402 (setton24; williams23), the “Rosetta” AGN (juodzbalis24a), and RUBIES-UDS-144195 (hviding25) are among the lowest-redshift LRDs, at , 3.19, 2.26, and 3.404, respectively. All but the “Rosetta” source show significantly redder UV slopes (see discussion in § IV). We additionally include the three reported extreme Balmer break sources (“black hole stars”, BH⋆) from degraaff25, naidu25, and taylor25, which exhibit weak UV emission and are identified solely by their extremely red optical colors.
@@
While the addition of these six spectroscopic LRDs contributes negligibly to the overall photometric sample, and less than 1% to the spectroscopic one, they highlight the need for a more complete diagnostic criterion based on rest-frame continuum properties. Classification using rest-frame UV and optical slopes (kocevski24; hainline24) or UV-to-optical colors (barro24) offers a more robust way to organize and distinguish LRD subtypes. As we show in § IV, these six sources fall at the red end of the rest-frame color diagram but are otherwise consistent with the broader LRD population.
@@
III.2 Other compact, non-LRD sources
Compact quiescent galaxies represent another population that can overlap with and contaminate LRD selections. These sources share key observable features, red colors, compact morphologies, and, in some cases, broad emission lines (kocevski23b), but their intrinsic SEDs differ markedly from those of LRDs. Specifically, quiescent galaxies exhibit red UV slopes, and while their optical continua are steep near the Balmer break, they flatten rapidly at longer wavelengths, deviating from the approximate power-law behavior typical of LRDs.
@@ III.2 Other compact, non-LRD sources
| Field | Area | Phot. LRDs | Spec. LRDs | S/N | BLAGNs LRDs | LRDs BL |
| CEERS-EGS | 88.6 | 127 | 45 | 25 | 61% | 89% |
| PRIMER-UDS | 243.0 | 353 | 68 | 33 | 59% | 89% |
| JADES-GDS | 62.1 | 92 | 21 | 17 | 50% | 70% |
| JADES-GDN | 56.0 | 112 | 19 | 16 | 67% | 54% |
| UNCOVER | 43.8 | 45 | 14 | 12 | 92% | 65% |
| PRIMER-COSMOS | 138.9 | 207 | 23 | 15 | N/A | N/A |
| Total | 632.4 | 968 | 191 | 118 | - | - |
III.3 Spectroscopic samples of BLAGNs
@@
Overall, we find that the fraction of LRDs among samples of BLAGNs is relatively high, typically ranging from 60% up to 90% in UNCOVER. These values are higher than those reported in earlier studies using the same BLAGN samples, primarily due to (1) the adoption of a more flexible photometric selection that extends to bluer UV-to-optical colors (i.e., shallower optical slopes), and (2) the inclusion of a small number of spectroscopic LRDs at lower redshift that fall outside traditional color cuts. For example, taylor24 report an LRD fraction of 34% among BLAGN in CEERS-EGS and PRIMER-UDS, while hainline24 find a value of 30% in JADES after excluding three LRD-targeted sources. Indeed, the spectroscopic follow-up strategy plays a significant role in these differences. The 90% fraction in UNCOVER reflects a strong targeting bias, as many sources were selected specifically to follow up the LRDs identified in labbe23. Similarly, the RUBIES survey prioritized red NIRCam sources, which naturally favors LRD-like objects. As noted by greene23, both the equivalent width and FWHM of broad lines tend to be larger in the reddest sources (e.g., F277WF444W 1.6), making BLAGNs that are also LRDs easier to detect. hainline24 and zhang25 reach a similar conclusion, showing that LRDs dominate the high-luminosity end of the AGN population. Consequently, it is clear that larger, unbiased spectroscopic samples are needed to reliably assess the intrinsic fraction of LRDs among the full BLAGN population.
The fraction of LRDs with broad emission lines is also relatively high, typically around 50–90%. Importantly, these values represent lower limits, as nearly all LRDs show clear emission lines, and the identification of broad components is often limited by the depth and sensitivity of the observations.
IV Rest-frame UV Slopes and Optical Colors
@@
IV.1 Rest-frame color - slope diagram classification
Here, we analyze a subset of 118 LRDs (out of 191) with NIRSpec/PRISM spectroscopy and high S/N continuum in the UV. We require an average S/N per pixel in the range Å. Since LRDs are typically red, this threshold also implies higher S/N in the optical, ensuring sufficient signal for reliable SED modeling. We compute UV slopes by fitting a power law to the continuum between the Lyman and Balmer limits, using the convention . The slope is measured in rather than the traditional , where , to enable direct comparison with the rest-frame UV-to-optical color, [0.30.9 m] , defined using synthetic tophat filters of width 100 Å.
@@ IV.1 Rest-frame color - slope diagram classification
This trend poses a challenge for photometric selections, which are tuned to detect the combination of blue UV and red optical slopes. Instead, the reddest LRDs begin to resemble quiescent or dusty galaxies with uniformly red continua, making them harder to identify in standard color–color diagrams (as noted in Section III.1). Several LRDs with red UV slopes were thus identified only a posteriori from spectroscopy (wang24_lrd; juodzbalis24a; williams24). Yet once placed on the color–slope diagram, they fall well within the main LRD distribution, supporting their classification as true LRDs.
@@
IV.2 Curvature vs. power-law in the rest-optical SED
We use the detailed spectral coverage of the optical SED provided by NIRSpec/PRISM to examine the continuum shape of LRDs. Our goal is to quantify what fraction of the population shows significant curvature or breaks in the optical, which may point to stellar-dominated SEDs (e.g., williams24; pg24a), or more exotic BH⋆ or quasi-stellar components (e.g., inayoshi24; naidu25). These cases differ from the smoother power-law continua often associated with canonical obscured AGNs in broadband photometry (e.g., polletta06).
@@
IV.3 The variety of SED types in LRDs
@@
The reddest bin ([0.30.9 m] mag) contains LRDs with the strongest curvature, reddest UV slopes, and most pronounced continuum breaks, with an average Balmer break strength of 4.41, far beyond what standard stellar population models can reproduce. This bin includes the three extreme LRDs from degraaff25, naidu25, and taylor25 that motivated the black hole star (BH⋆) model, in which a dense gas cocoon surrounds an accreting black hole, causing the sharp decline. We further discuss this modeling in § V. In addition to their breaks, these LRDs exhibit sharp flattening toward the NIR, consistent with the weak MIRI colors reported for the reddest LRDs (e.g., degraaff25; setton25; ronayne25). The adjacent bin ( [0.30.9 m] mag) shows similar overall shapes, with somewhat weaker curvature, red UV slopes, and a lower average Balmer break amplitude of 2.35. This group includes some of the first spectroscopically confirmed LRDs with strong but less extreme Balmer breaks (kocevski24; labbe24b; wang24b). In both of the two reddest bins, LRDs tend to exhibit prominent, broad Balmer emission lines, with weaker or undetected metal lines. We discuss line ratios and equivalent widths in more detail in § IV.7.
The three bottom bins correspond to LRDs with [0.30.9 m] , approaching the median color of the photometric sample ([0.30.9 m] mag). These sources display a remarkably similar flat UV continuum, with , much bluer than in the redder bins. As [0.30.9 m] decreases from 3 to 1, the average Balmer break amplitude declines steadily to 1.57, 1.13, and 1.07, values more typical of normal blue star-forming galaxies. Similarly, SED curvature diminishes toward shallower, power-law–like optical slopes. In these bins, the composite SED closely resembles the canonical combination of blue plus red power laws, with an inflection near the Balmer limit. In parallel, the median spectra show a growing prominence of narrow metal lines, especially the coronal lines around the Balmer break, such as [O ii] and [Ne iii], while the relative strength of H appears to decline compared to the blended H +[O iii]5007 complex.
Notably, the median SEDs in the two bluest bins suggest that the inflection may occur redward of H∞, near the [O iii]5007 emission line. This is consistent with “Virgil” (iani24; rinaldi25), a blue LRD with a seemingly flat optical continuum that rises sharply beyond H . A shift in the inflection is quite relevant. One of the main arguments against two-component models (e.g., galaxy+AGN) has been the consistency of the inflection at H∞, which would favor a single-component origin (setton24). If the inflection point varies with color, it would instead favor a multi-component scenario for the peculiar SED of LRDs. We revisit this scenario in detail in § V.
IV.4 Breakers vs. Power-Laws: Diversity and Degeneracy in LRD SEDs
@@
The left panel of Figure 6 shows LRDs with pronounced curvature, following the trend between [0.30.9 m] color and curvature strength (; see color-coding in Figure 4). The reddest, highest-curvature sources exhibit the most extreme Balmer breaks: their SEDs rise steeply in the optical and then flatten sharply toward the NIR. Moving toward bluer LRDs, the break becomes less pronounced, the optical slope shallower, and the flattening shifts to longer wavelengths. A simple interpretation in the context of a two-component model, would be that the relative luminosity of the break-dominated component gradually decreases top-to-bottom, effectively hiding the steep slope and pushing the inflection point redward.
The right panel of Figure 6 shows power-law LRDs. While most are blue, a small fraction (10%) exhibit very red colors ([0.30.9 m] in Figure 4). These red power-law sources show a range of optical slopes, suggesting variation in the physical conditions shaping the continuum, such as dust attenuation or gas density in BH⋆ models. Notably, the reddest of these LRDs, JADES-GDN-68797 and MACSJ0647.7+7015 (one of the first reported LRDs; killi23), exhibit remarkably similar optical slopes, despite the latter being about 1 mag bluer. This also agrees with the apparent shift in the inflection point from the flat UV to the rising optical continuum in this object, which occurs at longer wavelengths. A similar shift is seen in “Virgil” (yellow) and other bluer LRDs, as discussed in the previous section. We return to this point in § V.
IV.5 Color Dependence of the Emission Line FWHMs